Use the information below for Finch Company to answer the question that follow.

Finch Company began its operations on March 31 of the current year. Finch has the following projected costs:

April May June
Manufacturing costs (1) $159,000 $197,600 $201,300
Insurance expense (2) 980 980 980
Depreciation expense 2,100 2,100 2,100
Property tax expense (3) 590 590 590


(1) Of the manufacturing costs, three-fourths are paid for in the month they are incurred and one-fourth is paid for in the following month.
(2) Insurance expense is $980 a month; however, the insurance is paid four times yearly, in the first month of the quarter (i.e., January, April, July, and October).
(3) Property tax is paid once a year in November.

The cash payments expected for Finch Company in the month of April are

a.$159,000

b.$119,250

c.$122,190

d.$140,595
Answer #1

Answer:

Expected Cash Payment = Cash Payment for Manufacturing cost + Cash payment for Insurance

Cash Paid for Manufacturing cost in the month of April = 3/4 of April's Manufacturing Cost + 1/4 of March's Manufacturing Cost

Cash paid for Manufacturing Cost in the month of April = (3/4 * $159,000) + $0 = $119,250

Cash paid for Insurance = $980 * 3 = $2,940

Expected Cash payment in April = $119,250 + $2,940

Expected Cash Payment in April = $122,190